from a group of 5 men and 7 women, a committee of 2men and 3 women is to be formed. how many committees can be formed if 2 women are feuding and refuse to serve the community together?
Answer by sudhanshu_kmr(1152)


no. of ways to selecting 2 men and 3 women = 5c2 * 7c3 = 10 * 35 = 350

no. of ways when both 2 women are together in group = 5C2 * 5C1 = 10*5 = 50
( two women are already selected, we have to select only one woman from 5)


number of group in which both are not together = 350 - 50 = 300